# Emergence in graphs with near-extreme constraints

### v2

@@ 4.2. Variational equations
we get that $H^{\prime}(g(x,y))$ is $\Omega(\beta)$ on the diagonal rectangles that do not involve $I_{n+3}$ and is $-\Omega(\beta)$ on the off-diagonal blocks that do not involve $I_{n+3}$, with the exception of $I_{n+1}\times I_{n+2}$ and $I_{n+2}\times I_{n+1}$. This implies that $g(x,y)$ is exponentially small on the diagonal blocks, exponentially close to 1 on all but two of the off-diagonal blocks, and of course is close to $p$ on $I_{n+1}\times I_{n_{2}}$ and $I_{n+2}\times I_{n+1}$.
@@ 4.3. Maximizing worth
Let $C:[0,1]\to[0,1]$ be a function whose worth we aim to maximize. For each $i=1,2,\ldots,n+2$, let $a_{n+2}$ be the average values of $C(y)$ for $y\in I_{i}$. Of the terms contributing to $W(C)$, the entropy term is bounded above by $c\sum_{i=1}^{n}H(a_{i})+\frac{1-nc}{2}(H(a_{n+1})+H(a_{n+2}))$, since fluctuations in $C$ within each pode can only decrease the entropy. The edge density term is $-\alpha(c\sum_{i=1}^{n}a_{i}+\frac{1-nc}{2}(a_{n+1}+a_{n+2}))$.
@@
If we are at a maximum of $W$, then the gradient of $W$ must be zero and the Hessian must be negative semi-definite. The Hessian of $W$ with respect to the variables $\{a_{i}\}$ is precisely $-\beta M$ plus diagonal terms proportional to $H^{\prime\prime}(a_{i})$. The matrix $M$ is (nearly) zero on the diagonal, with all of the off-diagonal terms being close to 1 or $p$, and so has eigenvalues of both signs. The only way for the Hessian to be negative semi-definite is for all but one of entries $H^{\prime\prime}(a_{i})$ to be at least of order $\beta$. In other words, all columns that maximize $W$ must have every entry but one (or every entry) approximately equal to 0 or 1. In terms of $G$, for $y$ in any pode but one, $|G(x,y)-nc|$ must exceed $\Theta(1/\beta)$.
@@
  • If $G(x,y)\approx nc$ for $y\in I_{1}$, then $c\sum_{i=2}^{n}a_{i}+\frac{1-nc}{2}(a_{n+1}+a+{n+2})\approx nc$. But that is impossible if each $a_{i}$ (other than $a_{1}$) is equal to 0 or 1. Contradiction. Likewise, it is not possible to have $G(x,y)\approx nc$ for $y\in I_{2},\ldots,I_{n}$. The first $n$ variables $a_{i}$ are all either pegged to 0 or to 1.
@@ 4.4. Exact multipodality
(86) $-H^{\prime\prime}(g_{ij,0})\Delta g_{ij}\leq 2\beta\sum_{k=1}^{n+2}c_{k}(g_{ik}\Delta g_{jk}+g_{jk}\Delta g_{ik}),$
@@ 4.6. Distinct phases and rank
Each of the phases above the scallops proven in Theorems 6 and 7 have unique optimal graphons with distinct symmetries and cannot be analytically continued to one another.
@@
The optimal graphons described by Theorem 6 have rank 2, while the optimal graphons above the $n$-th scallop described by Theorem 7 have rank $n+2$. We will construct a sequence of “order parameters,” each a polynomial in finitely many subgraph densities, to distinguish between graphons of different rank. Specifically, the $k$th order parameter is identically zero whenever the rank of the optimal graphon is $k-1$ or less, and is never zero when the rank of the graphon is $k$. Since an analytic function on a connected set that is zero on an open subset is zero everywhere, there cannot be an analytic path connecting the $(k-2)$-nd scallop (where the graphon has rank $k$ and the order parameter is nonzero) to the previous scallops or to the $A(2,0)$ phase, where the order parameter is zero. In other words, the phases above the different scallops are all distinct.
@@
In summary: the $k$th order parameter is an analytic function of $(e,t)$ in each phase, being built from subgraph densities. It is identically zero on the regions above the 0th, 1st, $\ldots$, $(k-3)$rd scallops but is never zero on the region above the $(k-2)$nd scallop. Thus the region above the $(k-2)$nd scallop is in a different phase from the regions above all the previous scallops. Each scallop has its own unique phase.
@@ 5. Proof of Theorem 4
For each fixed $e\in(0,1)$ and all $t$ sufficiently close to (but below) $e^{3/2}$, the optimal graphon with edge/triangle densities $(e,t)$ is unique and bipodal. Asymptotically, the Boltzmann entropy scales as $-(e^{3/2}-t)\ln(e^{3/2}-t)$ and the Lagrange multiplier $\beta$ scales as $\ln(e^{3/2}-t)$.
@@
We follow the same overall roadmap as the proofs of Theorems 6 and 7. Specifically,
@@
Step 1 is identical to what we have done before. There is a unique graphon at $(e,e^{3/2})$, namely a graphon $g_{0}$ that is 1 on $I_{1}\times I_{1}$ and zero elsewhere, where $I_{1}$ is a pode of size $\sqrt{e}$. Every graphon with $t$ close to $e^{3/2}$, and in particular any entropy-maximizing graphon, must be $L^{2}$ close to $g_{0}$. This means that for all $x$’s outside of a set of small measure, $g_{x}$ is $L^{2}$-close to the corresponding column of $g_{0}$. This also implies that $G(x,y)$ is close to $\sqrt{e}$ when $x$ and $y$ are both in $I_{1}$ and is close to zero when either is in $I_{2}$.
@@
(101) $-\alpha\sqrt{e}-\frac{\beta e^{3/2}}{2}.$
@@
(102) $W(C)\leq\sqrt{e}H(a)+(1-\sqrt{e})H(b)-\alpha\sqrt{e}a-\alpha(1-\sqrt{e})b-\frac{\beta}{2}e^{3/2}a^{2},$
@@
Since $G(x,y)\approx 0$ or $\sqrt{e}$, depending on which quadrant we are in, this implies that $g(x,y)$ is exponentially close to 1 on $I_{1}\times I_{1}$ and exponentially close to 0 on $I_{1}\times I_{2}$ and $I_{2}\times I_{2}$. In particular, $H^{\prime\prime}(g)$, which scales as the larger of $1/g$ and $1/(1-g)$, is much larger than $|\beta|$. Looking at the change in the left hand side and right hand side of this equation within a single quadrant, we see that $H^{\prime\prime}(g)$ times the maximum fluctuation within any quadrant is of the same order as $|\beta|$ times the maximum fluctuation within any quadrant. But that means that the maximum fluctuation is bounded by a small multiple of itself, and so must be zero. Our graphon is exactly bipodal.
Finally, we do some calculations in the space of bipodal graphons. Let $g_{11}$, $g_{12}$ and $g_{22}$ be the values of the optimal graphon on $I_{1}\times I_{1}$, $I_{1}\times I_{2}$ and $I_{2}\times I_{2}$, respectively. We treat $g_{11}$, $g_{12}$ and $g_{22}$ as free variables and adjust the width of $I_{1}$ to keep the edge density fixed. To leading order, $e^{3/2}-t$ is a linear function of $(1-g_{11},g_{12},g_{22})$. However, $1-g_{11}$, $g_{12}$ and $g_{22}$ all scale as exponents of $\beta$, so $\beta$ must scale as $\ln(e^{3/2}-t)$. The entropy goes as $-(1-g_{11})\ln(1-g_{11})-g_{12}\ln(g_{12})-g_{22}\ln(g_{22})$, which then scales as $-(e^{3/2}-t))\ln(e^{3/2}-t)$.
@@
6. Inequivalence of ensembles
Using Lagrange multipliers is superficially similar to studying an exponential random graph model (ERGM; see Section 6 in [10], or [14], or [12] for a relevant introduction), where one starts with values $\alpha$ and $\beta$ and defines a free energy
(105) $\Psi(\alpha,\beta)=\max_{g}\left(S(g)-\alpha\varepsilon(g)-\frac{\beta}{3}\tau(g)\right).$
(In the literature, ERGMs are usually described in terms of parameters $\beta_{1}=-\alpha$ and $\beta_{2}=-\beta/3$, but that linear rescaling is not important.) The function $\Psi(\alpha,\beta)$ is the Legendre transform of the Boltzmann entropy $\mathbb{B}(e,t)$:
@@ 6. Inequivalence of ensembles
An LDP relates the graphon that maximizes the right hand side of (105), in other words that maximizes $F$ for the given values of $\alpha$ and $\beta$, to typical large graphs in a certain ensemble. In that alternate ensemble, all graphs on $n$ vertices are allowed, only with probability proportional to $\exp(-n^{2}(\alpha\varepsilon+\frac{\beta}{3}\tau))$, where $\varepsilon$ and $\tau$ are the edge and triangle densities of the graph in question.
If the Boltzmann entropy function were convex and the Razborov triangle were convex, then the Legendre transform (106) would be invertible within each phase [46]. If that were true, we could tune $\alpha$ and $\beta$ within each phase to get whatever values of $(e,t)$ we wanted, much as we can tune the temperature and pressure of a physical material to get whatever energy density and particle density that we want. In statistical mechanics, this ability to switch back and forth between fundamental variables and conjugate variables is called equivalence of ensembles. (Note that the equivalence of ensembles only involves invertibility in the interior of phases, not on the boundaries. The same temperature and pressure can describe a variety of water/ice mixtures with a range of energies.)
@@
It has long been known that all points with $t$ greater than $e^{3}$, and even moderately less than $e^{3}$, are ERGM-invisible [12]. The only points off the Erdős-Rényi curve that might be ERGM-visible lie close to the lower boundary of the triangle. We now show that, because of the nonconvexity of the Razborov triangle, most of those are also ERGM-invisible.
@@
The statement of Theorem 11 is not new; see Theorem 6.2 in [12]. However, the simplicity of the proof gets to the heart of why points in this region are ERGM-invisible.
Considering the Razborov triangle as a whole, only a few pieces are ERGM-visible. The Erdős-Rényi curve $t=e^{3}$ is ERGM-visible. Since $d\varepsilon$ and $d\tau$ are collinear at constant graphons, each point on the Erdős-Rényi curve actually corresponds to an infinite set of $(\alpha,\beta)$ values. A neighborhood of each cusp is ERGM-visible; that’s what you get when $\beta$ is large and positive. As far as we can tell, a neighborhood of the flat part of the lower boundary is also ERGM-visible. But that’s all. It appears that ERGMs can be used to study Erdős-Rényi graphs and parts of the $A(n,0)$ phases, but not the $F(1,1)$, $B(n,1)$ or $C(n,2)$ phases of Figure 2.

### v3

@@ 4.2. Variational equations
we get that $H^{\prime}(g(x,y))$ is $\Omega(\beta)$ on the diagonal rectangles that do not involve $I_{n+3}$ and is $\Omega(\beta)$ on the off-diagonal blocks that do not involve $I_{n+3}$, with the exception of $I_{n+1}\times I_{n+2}$ and $I_{n+2}\times I_{n+1}$. This implies that $g(x,y)$ is exponentially small (that is, $\exp(-\Omega(\beta))$) on the diagonal blocks, exponentially close to 1 on all but two of the off-diagonal blocks, and of course is close to $p$ on $I_{n+1}\times I_{n_{2}}$ and $I_{n+2}\times I_{n+1}$.
@@ 4.3. Maximizing worth
Let $C:[0,1]\to[0,1]$ be a function whose worth we aim to maximize. For each $i=1,2,\ldots,n+2$, let $a_{n+2}$ be the average values of $C(y)$ for $y\in I_{i}$. Of the terms contributing to $W(C)$, the entropy term is bounded above by $2c\sum_{i=1}^{n}H(a_{i})+(1-nc)(H(a_{n+1})+H(a_{n+2}))$, since fluctuations in $C$ within each pode can only decrease the entropy. The edge density term is $\alpha(2c\sum_{i=1}^{n}a_{i}+(1-nc)(a_{n+1}+a_{n+2}))$.
@@
If we are at a maximum of $W$, then the gradient of $W$ must be zero and the Hessian must be negative semi-definite. The Hessian of $W$ with respect to the variables $\{a_{i}\}$ is precisely $2\beta M$ plus diagonal terms proportional to $H^{\prime\prime}(a_{i})$. The matrix $M$ is (nearly) zero on the diagonal, with all of the off-diagonal terms being close to 1 or $p$, and so has eigenvalues of both signs. The only way for the Hessian to be negative semi-definite is for all but one of entries $H^{\prime\prime}(a_{i})$ to be at least of order $\beta$. In other words, all columns that maximize $W$ must have every entry but one (or every entry) approximately equal to 0 or 1. In terms of $G$, for $y$ in any pode but one, |G(x,y)nc||G(x,y)-nc| must exceed $\Theta(1/\beta)$.
@@
  • If $G(x,y)\approx nc$ for $y\in I_{1}$, then $2c\sum_{i=2}^{n}a_{i}+(1-nc)(a_{n+1}+a+{n+2})\approx 2nc$. But that is impossible if each $a_{i}$ (other than $a_{1}$) is equal to 0 or 1. Contradiction. Likewise, it is not possible to have $G(x,y)\approx nc$ for $y\in I_{2},\ldots,I_{n}$. The first $n$ variables $a_{i}$ are all either pegged to 0 or to 1.
@@ 4.4. Exact multipodality
(104) $H^{\prime\prime}(g_{ij,0})\Delta g_{ij}\leq 2\beta\sum_{k=1}^{n+2}c_{k}(g_{ik}\Delta g_{jk}+g_{jk}\Delta g_{ik}),$
@@ 4.5. Graphons with $(n,2)$ symmetry
As with Theorem 14, we have actually proven that an optimal graphon in $\mathcal{W}_{e,t}$ has a unique form up to group equivalence, implying
Corollary 18.
For all $(e,t)$ such that Theorem 17 applies, any graphon $g\in\mathcal{W}_{e,t}$ that is weakly equivalent to the multipodal entropy maximizer described in that theorem is actually group equivalent to the multipodal entropy maximizer.
@@ 4.6. Distinct phases and rank
Each of the phases above the scallops proven in Theorems 14 and 17 have unique optimal graphons with distinct symmetries and cannot be analytically continued to one another.
@@
The optimal graphons described by Theorem 14 have rank 2, while the optimal graphons above the $n$-th scallop described by Theorem 17 have rank $n+2$. We will construct a sequence of “order parameters,” each a polynomial in finitely many subgraph densities, to distinguish between graphons of different rank. Specifically, the $k$th order parameter is identically zero whenever the rank of the optimal graphon is k1k-1 or less, and is never zero when the rank of the graphon is $k$. Since an analytic function on a connected set that is zero on an open subset is zero everywhere, there cannot be an analytic path connecting the (k2)(k-2)-nd scallop (where the graphon has rank $k$ and the order parameter is nonzero) to the previous scallops or to the $A(2,0)$ phase, where the order parameter is zero. In other words, the phases above the different scallops are all distinct.
@@
In summary: the $k$th order parameter is an analytic function of $(e,t)$ in each phase, being built from subgraph densities. It is identically zero on the regions above the 0th, 1st, $\ldots$, (k3)(k-3)rd scallops but is never zero on the region above the (k2)(k-2)nd scallop. Thus the region above the (k2)(k-2)nd scallop is in a different phase from the regions above all the previous scallops. Each scallop has its own unique phase.
@@ 5. Proof of Theorem 4
For each fixed $e\in(0,1)$ and all $t$ sufficiently close to (but below) $e^{3/2}$, the optimal graphon with edge/triangle densities $(e,t)$ is unique and bipodal. Asymptotically, the Boltzmann entropy scales as $(e^{3/2}-t)\ln(e^{3/2}-t)$ and the Lagrange multiplier $\beta$ scales as $\ln(e^{3/2}-t)$.
@@
We follow the same overall roadmap as the proofs of Theorems 14 and 17. Specifically,
@@
  5. (5)
    Using Theorem 13 to eliminate the possibility that some points $(e,t)$ might have entropy maximizers without well-defined Lagrange multipliers $(\alpha,\beta)$.
Step 1 is identical to what we have done before. There is a unique reduced graphon at $(e,e^{3/2})$, namely the equivalence class of a graphon $g_{0}$ that is 1 on $I_{1}\times I_{1}$ and zero elsewhere, where $I_{1}$ is a pode of size $\sqrt{e}$. Every graphon with $t$ close to $e^{3/2}$, and in particular any entropy-maximizing graphon, must be $L^{2}$ close to $g_{0}$. This means that for all $x$’s outside of a set of small measure, $g_{x}$ is $L^{2}$-close to the corresponding column of $g_{0}$. This also implies that $G(x,y)$ is close to $\sqrt{e}$ when $x$ and $y$ are both in $I_{1}$ and is close to zero when either is in $I_{2}$.
@@
(119) $2\alpha\sqrt{e}-\beta e^{3/2}.$
@@
(120) $W(C)\leq 2\sqrt{e}H(a)+2(1-\sqrt{e})H(b)-2\alpha\sqrt{e}a-2\alpha(1-\sqrt{e})b-\beta e^{3/2}a^{2},$
@@
Since $G(x,y)\approx 0$ or $\sqrt{e}$, depending on which quadrant we are in, this implies that $g(x,y)$ is exponentially close to 1 on $I_{1}\times I_{1}$ and exponentially close to 0 on $I_{1}\times I_{2}$ and $I_{2}\times I_{2}$. In particular, $H^{\prime\prime}(g)$, which scales as the larger of $1/g$ and 1/(1g)1/(1-g), is much larger than $|\beta|$. Looking at the change in the left hand side and right hand side of this equation within a single quadrant, we see that $H^{\prime\prime}(g)$ times the maximum fluctuation within any quadrant is of the same order as $|\beta|$ times the maximum fluctuation within any quadrant. But that means that the maximum fluctuation is bounded by a small multiple of itself, and so must be zero. Our graphon is exactly bipodal.
Finally, we do some calculations in the space of bipodal graphons. Let $g_{11}$, $g_{12}$ and $g_{22}$ be the values of the optimal graphon on $I_{1}\times I_{1}$, $I_{1}\times I_{2}$ and $I_{2}\times I_{2}$, respectively. We treat $g_{11}$, $g_{12}$ and $g_{22}$ as free variables and adjust the width of $I_{1}$ to keep the edge density fixed. To leading order, $e^{3/2}-t$ is a linear function of $(1-g_{11},g_{12},g_{22})$. However, $1-g_{11}$, $g_{12}$ and $g_{22}$ all scale as exponents of $\beta$, so $\beta$ must scale as $\ln(e^{3/2}-t)$. The entropy goes as $(1-g_{11})\ln(1-g_{11})-g_{12}\ln(g_{12})-g_{22}\ln(g_{22})$, which then scales as $(e^{3/2}-t))\ln(e^{3/2}-t)$.
@@
Finally, as with Theorems 14 and 17, we note that we have determined the entropy maximizer up to group equivalence, implying
Corollary 21.
Let $(e,t)$ be such that Theorem 20 applies. Then any graphon $g\in\mathcal{W}_{e,t}$ that is weakly equivalent to the bipodal maximizer defined in the theorem is actually group equivalent to that bipodal maximizer.
6. ERGM-invisibility
Using Lagrange multipliers is superficially similar to studying an exponential random graph model (ERGM; see Section 6 in [11], or [15], or [13] for a relevant introduction), where one considers an ensemble of all graphs on $n$ vertices, where the probability of a given graph $G$ is proportional to
(123) $\exp\left(-n^{2}\left[\alpha\varepsilon(G)+\frac{\beta}{3}\tau(G)\right]\right).$
Here $\varepsilon(G)$ and $\tau(G)$ are the edge and triangle densities of the graph $G$ and $\alpha$ and $\beta$ are variables which can move the distribution, expected to be narrowly peaked for large $n$. (In the literature, ERGMs are usually described in terms of parameters $\beta_{1}=-\alpha$ and $\beta_{2}=-\beta/3$, but that linear change of variables does not matter.)
ERGMs are widely used to model real-world networks, with $n$ necessarily finite. See [20], [21] and their bibliographies for references relevant to the current discussion. By 2010 there was literature noting that fitting of parameters to data had various problems, and in [13] the recent LDP for $\mathbb{G}(n,p)$ graphs [14] was applied to see if that would help. See [9] and the introduction in [13] for background. [13] illuminated some issues but left some others unresolved. Our treatment of the Boltzmann entropy can help, as follows.
The $n\to\infty$ limit of an ERGM with parameters $(\alpha,\beta)$ can be understood in terms of graphons and the function
(124) $\Psi(\alpha,\beta)=\max_{g}\left(S(g)-\alpha\varepsilon(g)-\frac{\beta}{3}\tau(g)\right)\equiv\max_{g}F(g).$
The LDP relates the graphon that maximizes $F(g)$ on the right hand side of (124), for the given values of $\alpha$ and $\beta$, to typical large graphs in the ensemble. Note that $\Psi(\alpha,\beta)$ is the Legendre transform of the Boltzmann entropy $\mathbb{B}(e,t)$:
@@ 6. ERGM-invisibility
If the Boltzmann entropy function were convex and the Razborov triangle were convex, then the Legendre transform (125) would be invertible within each phase [52]. If that were true, we could tune $\alpha$ and $\beta$ within each phase to get whatever values of $(e,t)$ we wanted. In statistical mechanics, this ability to switch back and forth between fundamental variables and conjugate variables is called equivalence of ensembles.
@@
It has long been known that all points with $t$ greater than $e^{3}$, and even moderately less than $e^{3}$, are ERGM-invisible [13]. The only points off the Erdős-Rényi curve that might be ERGM-visible lie close to the lower boundary of the triangle (Figure 1). We now show that, because of the nonconvexity of the Razborov triangle, most of those are also ERGM-invisible.
@@
The statement of Theorem 23 is not new; see [13, Theorem 6.2]. However, the simplicity of the proof gets to the heart of why points in this region are ERGM-invisible.
Considering the Razborov triangle as a whole, only a few pieces are ERGM-visible. The Erdős-Rényi curve $t=e^{3}$ is ERGM-visible. Since $d\varepsilon$ and $d\tau$ are collinear at constant graphons, each point on the Erdős-Rényi curve actually corresponds to an infinite set of $(\alpha,\beta)$ values. A neighborhood of each cusp is ERGM-visible; that’s what you get when $\beta$ is large and positive. As far as we can tell, a neighborhood of the flat part of the lower boundary is also ERGM-visible. In short, ERGMs can see homogeneous Erdős-Rényi or $A(n,0)$ structures, where every vertex looks like every other vertex. But that’s all.
In this paper, we have shown how a homogeneous constraint on the total number of edges and triangles leads to inhomogeneous $F(1,1)$ or $C(n,2)$ structures, where vertices in one set of podes look very different from vertices in another set of podes. We previously showed similar behavior near the Erdős-Rényi curve. However, ERGMs cannot see this spontaneous emergence of inhomogeneity, as the portions of the Razborov triangle where such emergent inhomogeneity occurs are all ERGM-invisible.
This is not to deny the real successes that ERGMs have had. Starting with [14] and [13], definitions were given of phases in the $(\alpha,\beta)$ plane. The existence of phases and phase transitions was proven quickly, far faster than for the analogous problems discussed in this paper. Those are important triumphs. Fundamentally, the two approaches to studying large graphs represent different parts of applied mathematics, with different goals.